Hi, I was wondering if someone could help me. I have no idea how to complete this problem, and I need to show my work (steps). So, if anyone has any help they can provide, I would greatly appreciate it. Thank you.
Simplify

cubrt(27x^4y^10).

Explain your answer and show all steps.

Answer by sudhanshu_kmr(1152)


cubrt(27 x^4 y^10 )

= (27 x^4 y^10)^(1/3)

= (3^3 x^4 y^10)^(1/3)

= 3 ^(3 * 1/3 ) x^(4* 1/3) y^(10 * 1/3)

= 3^1 x^(4/3) y^(10/3)

= 3 x^(4/3) y^(10/3)